OpenStudy (anonymous):

A canoeing club went on a 9-day trip. The first day, the club canoed 4 2/3 miles. Each day after that, the club canoed between 9 2/3 and 11 3/4 miles. Which could be the total number of miles the club canoed by the end of the trip? A. 80 B. 90 C. 101 D. 117

OpenStudy (anonymous):

the min they canoed is 4 2/3 + 8(9 2/3)=82 mi the max they canoed is 4 2/3 + 8(11 3/4)=98 2/3 mi

OpenStudy (anonymous):

so, option b is the only viable choice

OpenStudy (anonymous):

Yep, if you estimate 8 * 10 = 80 80 + 4 = 84 => only choice B = 90 is closest one!